OpenStudy (anonymous):

Medals shall be given to the best response :) Please help, I have difficulty solving such questions. A photocopying store charges a flat rate of $2 plus $0.05/copy. (a) Write a function f(x) to represent the average cost per copy. (b) Determine what happens to the function as x becomes very large. (c) What is the significance of this value? Thanks for your help.

OpenStudy (anonymous):

I forgot to add this but no calculus methods should be used in solving this problem.

OpenStudy (ivettef365):

you have a flat fee of 2, and each copy .05 so for number of copies use x (a variable for unknown) so you have f(x) = 2 + .05x

OpenStudy (anonymous):

Let x be number of copies. Total Cost for Photocopying X = Flat Cost + Pro-rated Cost = 2 + (0.05)(x) = f(x) Average Cost for Photocopying X = (2+0.05x) / x Expanding: 2/x + 0.05 When x tends towards infinity (becomes very large), 2/x tends towards 0, thus Average Cost tends towards 0.05. Significance of this value is that average cost for photocopying will tend towards the variable component.

OpenStudy (anonymous):

Thank you all so much. I wish I could give all of you medals but it goes to the best response. Thank you so much.